OpenStudy (dj3strella):

Help Factor x^2 – 4x + 4

OpenStudy (anonymous):

seriously? use competing square method (x)^2 -2(x)(2) + (2)^2 -(2)^2 +4 =0 (x-2)^2=0 x=2

OpenStudy (dj3strella):

thank you but that's not what i was looking for

OpenStudy (dj3strella):

i no longer need your help. you have failed me

OpenStudy (anonymous):

its the right answer!

OpenStudy (anonymous):

factors are (x-2)(x-2)